Perreth Products manufactures its products in two separate departments: Machining and Assembly. Total manufacturing overheadMachining Assembly Department Department Job 500 ... 10 MH 12 DL hours 4 DL hours Job 501 20 MH 12 DL hours 4 DL hours Both JRequirement 1. Compute the companys current plantwide overhead rate. (Round your answer to the nearest dollar.) Begin by detRequirement 4. Compute the total amount of overhead allocated to each job if the company uses its current plantwide overheadRequirement 7. Compute the total manufacturing cost and sales price of each job using the companys current plantwide overhea
Answer #1

1.Predetermined Plantwide rate = Total Overheads/Total labor hours

=1050000/15,000 = $70 per labor hour

2.Rate = Departmental Overheads/Driver

Machining = 630,000/8400 = $75 per machine hour

Department B = 420,000/12,000 = $35 per labor hour

501 uses more

10 more machine hours, costs the company more resources

Job 500

Job 501

Total direct labor hours

16

16

Plantwide

70

70

Overhead

1,120

1,120

Job 500

Job 501

Machining

=75*10 = $750

1,500

Assembly

420

420

Total

1,170

1,920

Equal overheads

More to 501

extra machine hours used

Correct/fair

Job 500

Job 501

Direct Material

1500

1500

Direct labor

480

480

Manufacturing overhead

1,120

1,120

Total manufacturing costs

3100

3100

Mark up

125%

125%

Sales Price

3,875

3,875

Job 500

Job 501

Sales Price

3,875

3,875

Total manufacturing costs

3,100

3,100

Gross Profit

775

775

Job 500

Job 501

Sales Price

3,875

3,875

Less: Costs

Direct Material

1500

1500

Direct labor

480

480

Manufacturing overhead

1170

3150

1920

3,900

Gross Profit

725

-25

Believes that it earns

Realises that loses money on 501